OpenStudy (anonymous):

pLEASE CHECK MY WORK TO SEE IF i AM CORRECT 28x^2-6x+1 28(-4)^2-6(4)+1 Then I need help

OpenStudy (anonymous):

(gof) (-4)

OpenStudy (amistre64):

you need to use -4 in each x spot ...

OpenStudy (anonymous):

I thought that I did that above 28(-4)-6(-4)+1 right

OpenStudy (anonymous):

sorry for got the 28(-4)^2

OpenStudy (amistre64):

better, and remember that -*- = +

OpenStudy (anonymous):

so that would be 28*4=112^2 right

OpenStudy (amistre64):

28(-4)^2 -6(-4) + 1 28(16) +24 + 1 280 + 120 + 48 +25

OpenStudy (amistre64):

28*4 = 80 + 32 = 112 yes

OpenStudy (anonymous):

so 112-24+1=89 but my test said that was wrong so would you add 112+24+1

OpenStudy (amistre64):

the 112 still needs to be multiplied by the other 4

OpenStudy (amistre64):

112*4 = 400+40+8

OpenStudy (anonymous):

so 448

OpenStudy (amistre64):

and its not -24 .. -6(-4) = 24, to negatives gets a postive

OpenStudy (anonymous):

that is right two negatives make positive so the answer would be 448 correct

OpenStudy (amistre64):

448 + 24 + 1 = ?

OpenStudy (anonymous):

the answer was 473???

OpenStudy (anonymous):

Ok I have four problems that goes with the pair of functions can you help me and I wont bug you no more tonight

OpenStudy (amistre64):

448 + 24 + 1 468 + 4 + 1 472 + 1 = 473

OpenStudy (amistre64):

i cant make that promise, its getting late over here where i am and il have to leave soon

OpenStudy (anonymous):

f(x)=3x g(x)=7x^2-4x+6

OpenStudy (amistre64):

post it up as a new question and see if anyone bites :)